OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form !!!!

OpenStudy (anonymous):

use this equation:\[y- y_1 = m(x-x_1)\] \[y_1 \text{ and } x_1\] are the x and y coordinates of a point on the line and m is the gradient \[y - 4 = 2(x - (-3))\] now you can rearrange to whatever form you like :)

OpenStudy (anonymous):

would it be y-10=2x ?

OpenStudy (anonymous):

kidding !